OpenStudy (anonymous):

Find, to the nearest degree, the acute angle of rotation that eliminates the xy term of -3x^2+3xy+5y^2-x+2y-6=0. a.)81° b.)41° c.)40° d.)80°

OpenStudy (superdavesuper):

the curve looks like this but no idea how to eliminate the xy term :(

OpenStudy (ranga):

If the general equation is: Ax^2 + Bxy + Cy^2 + Dx + Ey + F = 0 you can eliminate the Bxy term by rotating an angle of theta where tan(2*theta) = B / (A - C) Here, B = 3, A = -3, C = 5. tan(2*theta) = 3 / (-3 - 5) = -3/8 2*theta = arctan (-3/8) theta = 1/2 * arctan (-3/8)

OpenStudy (ranga):

arctan(-3/8) = -20.556 or +159.444 Choosing the positive angle, 2 * theta = 159.444 theta = 159.444 / 2 = 79.722 nearest degree is 80 degrees.

OpenStudy (superdavesuper):

Nice work @ranga!! I was stumped yesterday but I was staring at the plot while you typed up ur solution n noticed that if I rotate them to the "East-West opening hyperbola" form, the xy term will disappear. So an alternative way to find the angle is by solving for the slope of its transverse axis. It comes out to be 5.513 n arctan(5.513) gives the same 79.7 angle :)

OpenStudy (ranga):

Exactly @superdavesuper That is another way to do it. Well done!